Find the first three terms of the recursively defined sequence.
a1 = 5, a(n) = na(n−1)

Answer by greenestamps(13200)



The statement

a(n) = na(n-1)

means that, for each integer n greater than 1, the n-th term is n times the previous term.

Given a(1)=5, use that definition to find a(2); then use the definition again to find a(3).

Here is an example if you don't understand that....

a(1) = 4; a(n) = n*a(n-1)+1

The second statement says to get the n-th term of the sequence you multiply the current term by n and add 1.

given first term: 4
second term: multiply 1st term (4) by n (2) and add 1: 4(2)+1 = 9
third term: multiply second term (9) by n (3) and add 1: 9(3)+1 = 28
fourth term: 28(4)+1 = 113